the best paper was wrongly scored 75 instead of 85 what will be the new mean if the average score of 24 students in a class in 58 marks


1
Expert's answer
2021-03-22T08:43:19-0400

"\text{An incorrect old average score:}"

"\bar{x_{1}}=\frac{1}{n}\Sigma_{i=1}^nx_i;n=24"

"\bar{x_{1}}=\frac{1}{24}\Sigma_{i=1}^{24}x_i=58;"

"\text{The correct average score:}"

"\bar{x}=\frac{1}{24}(\Sigma_{i=1}^{24}x_i+10)=58+\frac{10}{24}\approx58.42"

Answer: 58.42 is a new average score.